# Enhancing Table Recognition with Vision LLMs: A Benchmark and Neighbor-Guided Toolchain Reasoner

### v2

@@ Abstract.
Pre-trained foundation models have recently significantly progressed in structured table understanding and reasoning. However, despite advancements in areas such as table semantic understanding and table question answering, recognizing the structure and content of unstructured tables using Vision Large Language Models (VLLMs) remains under-explored. In this work, we address this research gap by employing VLLMs in a training-free reasoning paradigm. First, we design a benchmark with various hierarchical dimensions relevant to table recognition. Subsequently, we conduct in-depth evaluations using pre-trained VLLMs, finding that low-quality image input is a significant bottleneck in the recognition process. Drawing inspiration from these findings, we propose the Neighbor-Guided Toolchain Reasoner (NGTR) framework, which is characterized by integrating multiple lightweight models for low-level visual processing operations aimed at mitigating issues with low-quality input images. Specifically, we utilize a neighbor retrieval mechanism to guide the generation of multiple tool invocation plans, transferring tool selection experiences from similar neighbors to the given input, thereby facilitating suitable tool selection. Additionally, we introduce a reflection module to supervise the tool invocation process. Extensive experiments on public table recognition datasets demonstrate that our approach significantly enhances the recognition capabilities of the vanilla VLLMs. We believe that the designed benchmark and the proposed NGTR framework could provide an alternative solution in table recognition [^1] Our code is open-sourced and available at https://github.com/lqzxt/NGTR.
Vision Large Language Models, Table Recognition, Tool Invocation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Tables are ubiquitous for organizing and communicating structured information across diverse domains, ranging from scientific literature and business reports to web pages and financial documents (zhang2020web, ; zheng2021global, ). They encapsulate a wealth of information essential for numerous applications, such as knowledge discovery, decision support, and data-driven analytics (shwartz2022tabular, ; wang2024tabletime, ). In the context of intelligent table applications, one fundamental yet challenging task is table recognition (zanibbi2004survey, ): converting image-based table representations into structured digital formats. Over the years, substantial efforts (salaheldin2024deep, ) have been made to tackle this problem, introducing various approaches to address challenges such as segmentation techniques and object detection methods.
@@
Figure 1. Comparison of modeling paradigms: domain-specific lightweight models vs. universal pre-trained VLLMs.
@@
Recently, the advent of Large Language Models (LLMs) (chang2024survey, ) and Vision Large Language Models (VLLMs) (yin2024survey, ) has revolutionized natural language processing and computer vision, respectively. For LLMs, their ability to process serialized data has facilitated numerous tabular data mining tasks, such as table-to-text generation (liu2018table, ; cheng2021hitab, ; guo2024adapting, ), table question answering (wang2024chain, ; gao2023text, ; mao2024potable, ), and table semantic understanding (deng2022turl, ; feuer2023archetype, ; trabelsi2020semantic, ). Meanwhile, several VLLM-based methods have emerged to bypass traditional OCR pipelines for visual table analysis and understanding (zheng2024multimodaltableunderstanding, ; hu2024mplug, ).
@@
Despite these advancements, our investigation reveals a noticeable gap: the application of VLLMs to table recognition remains underexplored. This task serves as a foundational building block for table-related applications. Some existing work (chen2023tablevlm, ; zhang2024multi, ) has focused on pre-training or fine-tuning VLLMs to accomplish this task. However, Fine-tuning VLLMs for specific tasks is often computationally expensive and risks catastrophic forgetting of general capabilities. To address this, we focus on leveraging a prompt-based paradigm for table recognition using pre-trained VLLMs, exploring a generative approach without requiring additional fine-tuning. Recognizing the absence of dedicated benchmarks in this domain, we design an evaluation framework based on a hierarchical design philosophy (sui2024table, ; cheng2023formertime, ; liu2024generative, ), comprising recognition tasks for table recognition. Through extensive evaluations, we identify a critical bottleneck: low-quality input images significantly hinder the table recognition capabilities of the evaluated VLLMs. This is understandable given the various characteristics associated with different table styles.
@@
To overcome this limitation, we propose the Neighbor-Guided Toolchain Reasoner (NGTR) framework with the goal of effective table recognition. One of the main characteristics of the NGTR framework is that it integrates the capacity of lightweight models and the strategy of retrieval-augmented generation to enhance input image quality and guide structured data recognition. Specifically, we propose a preprocessing toolkit equipped with various lightweight models designed to enhance input image quality. For each input instance, we retrieve a similar neighbor from the training data, and the experience gained from these neighbors is used to inform the guidance for the generation of tool invocation plans. This experience based guidance helps select and integrate suitable tools for the given input. Furthermore, we incorporate a reflection-driven tool selection module at each step of tool invocation to iteratively refine table recognition outputs. This enables VLLMs to produce higher-quality structured data.
@@
To validate the effectiveness of the proposed NGTR framework, we conduct extensive experiments on multiple public table recognition datasets. The key observations are as follows: (1) Our NGTR framework significantly enhances the table recognition performance of naive VLLM-based approaches; (2) While VLLMs achieve competitive accuracy on specific datasets compared to traditional models, a noticeable performance gap remains in favor of traditional models. Nonetheless, we have preliminarily revealed the performance boundaries of VLLMs in several representative table recognition datasets. As is shown in Figure 1, the VLLM-based table recognition approach demonstrates the capability for universal modeling. This method facilitates a paradigm shift in design objectives from a model-centric to a data-centric focus, presenting significant potential for further exploration. We hope this work will inspire more research efforts in the future.
In summary, the contributions of this paper are as follows:
  • We pioneer the exploration of VLLM-based table recognition tasks by introducing a comprehensive benchmark with hierarchical evaluation settings.
@@ 2. Related Work
Table Recognition. Earlier table recognition (TR) methods predominantly rely on heuristic rules (TR-Heu-1, ; TR-Heu-2, ; TR-Heu-3, ) or statistical learning techniques (TR-Stat-1, ; TR-Stat-2, ), which were generally effective on simple and standard structured tables. However, these approaches rely heavily on handcrafted features or implicit rules and show limited generalization ability. In the era of deep learning, numerous studies have made impressive progress in handling more intricate and heterogeneous table structures. Top-down methods (TR-TD-1, ; TR-TD-2, ; TR-TD-3, ; qin2024semv3, ) first predict table lines and boundaries to infer the overall structure, and then locate cells based on row-column intersections. Bottom-up methods (zheng2021global, ; qiao2021lgpma, ; xing2023lore, ) first identify table cells with object detection models (OD-1, ; OD-2, ), and then predict the cell relations to organize the row-column structures to form the overall tables. These methods follow an explicit two-stage learning paradigm with relatively strong transferability and explainability, yet the risks of ambiguous contents and boundless structures may cause unstable and incorrect predictions.
@@
Recently, sequence-based methods (zhong2020image, ; TR-Seq-2, ; TR-Seq-3, ; TR-Seq-4, ) have been widely explored to directly generate markup sequence defining structures with specific decoders. Although these approaches require a massive of training data and computing resources, they have demonstrated substantial potential to unify visual-text parsing tasks (wan2024omniparser, ).
@@
Large Language Models. In recent years, LLMs have made significant progress in the field of natural language processing. In particular, LLMs such as GPT (radford2018improving, ) and Llama (touvron2023llama, ) have demonstrated exceptional performance in tasks such as multi-task learning (chen2024multi, ), zero-shot learning (kojima2022large, ), and text generation (li2024pre, ). These models have not only broken through the limitations of traditional technologies in generating natural language text, but have also shown capabilities in reasoning (wei2022chain, ; cheng2024advancing, ; shinn2024reflexion, ) and planning (guan2023leveraging, ; gou2023critic, ). Meanwhile, VLLMs combine the capabilities of visual and language understanding, enabling LLMs to process visual information. For multimodal understanding scenarios (e.g., scene text recognition (STR, ), visual question answering (VQA, )), VLLMs have been widely validated as effective (Image2LLM, ; UReader, ; TextMonkey, ). With continuous progress in text-rich scenarios, some studies (zheng2024multimodaltableunderstanding, ; zhao2024tabpedia, ; chen2023tablevlm, ) have also focused on enabling VLLMs to handle multimodal table understanding tasks.
@@
Despite their promising success in a wide range of domains, VLLMs applied to table recognition remain under-evaluated and under-explored. Our study presents a comprehensive benchmark for VLLM-based table recognition evaluation. Subsequently, we propose a novel framework to address the bottleneck of VLLMs, thereby enhancing their capabilities in table recognition.
3. Problem Definition and Proposed Benchmark
3.1. Problem Definition
@@
3.2. Benchmark Evaluation Setup
@@
Table 1. Descriptions of the proposed hierarchical recognition tasks.
@@
3.2.1. Recognition Task Design
Cell-level. We evaluate the cell-level recognition capability of VLLMs within three specific recognition tasks. Merged cell detection task aims to recognize cells that span multiple rows or columns in the table. Formally, given an image $I$, this task aims to recognize $\{c_{1},c_{2},\ldots,c_{n}\}$ that represents the textual content of all merged cells in the table. Content/index-based cell recognition tasks evaluate the structural and content recognition capabilities of VLLMs, which are crucial for assessing whether VLLMs could perform well in fine-grained table recognition. Formally, define an image $I$, the cell content $c$, and the cell’s row and column indices $(R,C)$. The Content-based cell recognition task identifies the logical location $(R,C)$ of the target cell, and the index-based cell recognition task extracts the content of the target cell $c_{(R,C)}$.
Row/Column-level. We evaluate the row/column-level recognition capability of VLLMs within two specific recognition tasks. Row/column index-based data recognition tasks are designed to assess whether VLLMs can accurately identify row/column elements. Formally, given a table image $I$ with specific cell’s row index $R$ or column index $C$, the VLLMs are required to return the entire row content $[c_{(R,1)},c_{(R,2)},\ldots]$ or the column content $[c_{(1,C)},c_{(2,C)},\ldots]$.
Table-level. We evaluate the overall table-level recognition capability of VLLMs. The visual table size detection recognition task is designed to evaluate whether VLLMs can accurately determine the number of rows and columns in a table, assessing their comprehension of the global structural information.
3.2.2. Baselines
@@
In this study, we evaluate the performance of six VLLMs. For open-source VLLMs, we select Phi-3.5-Vision-Instruct (Phi) and Llama-3.2-90B-Vision-Instruct (Llama) for evaluation. For closed-source VLLMs, we evaluat GPT-4o-mini (GPT-mini), Qwen-VL-Max (Qwen), GPT-4o (GPT) and Gemini-1.5-Pro (Gemini). The specific details of VLLM can be found in the appendix D.
@@
Figure 2. Experimental results of VLLMs for the proposed hierarchical tasks. The tasks evaluated include the following: merged cell detection (MCD), content/index-based cell recognition (CCR, ICR), index-based row/column data recognition (IRDR, ICDR), and visual table size detection (VTSD).
3.3. Benchmark Evaluation Results
@@
Comprehensive evaluations of hierarchical structural recognition tasks are presented in Figure 2. The results show that, among all the VLLMs we selected, GPT and Gemini demonstrate the strongest performance, consistently outperforming the other VLLMs. Furthermore, open-source Llama demonstrates a significant performance gap compared to the closed-source VLLMs. We give some highlights associated with the benchmark results as follows:
Figure 3. Row-column sensitivity analysis of VLLMs on designed hierarchical tasks. The left figure uses Gemini, and the right figure uses GPT.
@@
Row-column Sensitivity Analysis. We found that all models show inconsistent performance between row/column-related tasks. To mitigate the influence of uneven distributions of rows and columns in the original data leading to varying difficulty levels, we further refine the experimental results by selecting samples where the difference between the number of rows and columns does not exceed three. The refined results are presented in Figure 3. The results of row-column sensitivity analysis experiments show that the accuracy for the column-related tasks is higher than for the row-related tasks, suggesting that VLLMs prefer to process column-structured information.
We believe this phenomenon is because columns usually contain more substantial attribute information, allowing VLLMs to locate columns more accurately. Rows tend to have higher similarities, so VLLMs faces more significant challenges locating rows or completing row-related tasks. The differences between rows and columns in visual table size detection further verify this view.
@@
Image Quality Analysis. We conduct in-depth experiments and analysis on the TR task. Detailed experimental results and analysis are presented in Section 5. The results demonstrate that significant bottlenecks remain despite existing VLLMs achieving competitive performance on the TR task. Using simple prompts, the publicly available VLLMs performs relatively well on the SciTSR dataset with higher image quality, scoring only 4.38 TEDS points lower than the leading baseline models. However, when processing the PubTabNet dataset with lower image quality, the performance gap widens significantly, reaching 13.63 TEDS points. This phenomenon indicates that the quality of the input image is a key bottleneck limiting the performance of VLLMs. Section 3.4 provides a more in-depth analysis of this bottleneck.
@@
Table 2. Specific description of the challenges and scenarios in low-quality image inputs.
@@
3.4. Bottleneck Analysis
In this section, we further investigate the performance of VLLMs under varying image quality conditions and assess their visual robustness to these conditions through empirical analysis.
3.4.1. Experimental Setup
To comprehensively evaluate the visual robustness of VLLM, we focus on three distinct visual challenges: the image quality challenge, the table border quality challenge, and the geometric deformation challenge. Specifically, We randomly selected 1,000 samples from the SciTSR (chi2019complicated, ) test set for our experiments and configure our image types according to these challenges. Details of these challenges are provided in Table 2.

### v3

@@ Abstract
Pre-trained foundation models have recently made significant progress in table-related tasks such as table understanding and reasoning. However, recognizing the structure and content of unstructured tables using Vision Large Language Models (VLLMs) remains under-explored. To bridge this gap, we propose a benchmark based on a hierarchical design philosophy to evaluate the recognition capabilities of VLLMs in training-free scenarios. Through in-depth evaluations, we find that low-quality image input is a significant bottleneck in the recognition process. Drawing inspiration from this, we propose the Neighbor-Guided Toolchain Reasoner (NGTR) framework, which is characterized by integrating diverse lightweight tools for visual operations aimed at mitigating issues with low-quality images. Specifically, we transfer a tool selection experience from a similar neighbor to the input and design a reflection module to supervise the tool invocation process. Extensive experiments on public datasets demonstrate that our approach significantly enhances the recognition capabilities of the vanilla VLLMs. We believe that the benchmark and framework could provide an alternative solution to table recognition [^1] Our code is available at https://github.com/lqzxt/NGTR.
1 Introduction
@@
Tables are ubiquitous for organizing and communicating structured data across diverse domains, ranging from scientific literature and business reports to web pages and financial documents Ye et al. (2024); Zheng et al. (2021). They store a wealth of information essential for applications such as knowledge discovery, decision support, and data-driven analytics Shwartz-Ziv and Armon (2022); Wang et al. (2024a). In the context of intelligent table applications, one fundamental yet challenging task is table recognition Zanibbi et al. (2004): converting image-based table representations into structured data formats. Over the years, substantial efforts Salaheldin Kasem et al. (2024) have been made to address this problem, introducing various approaches to address challenges, such as image segmentation techniques and cell object detection methods.
@@
Figure 1: Comparison of modeling paradigms: domain-specific lightweight models vs. universal pre-trained VLLMs.
@@
Recently, the advent of Large Language Models (LLMs) Chang et al. (2024) and Vision Large Language Models (VLLMs) Yin et al. (2024) has revolutionized natural language processing and computer vision. For LLMs, their powerful understanding and reasoning capabilities have facilitated numerous tabular data mining tasks, such as table-to-text generation Guo et al. (2024), table question answering Wang et al. (2024b); Mao et al. (2024), and table semantic understanding Deng et al. (2022); Cheng et al. (2025a). Meanwhile, several VLLM-based methods have emerged to bypass traditional OCR pipelines for visual table analysis and understanding Hu et al. (2024).
@@
Despite these advancements, our investigation reveals a noticeable gap: the application of VLLMs to table recognition remains underexplored. This task serves as a foundational building block for table-related applications. Some existing work Luo et al. (2024); Zhang et al. (2024) has focused on pre-training or fine-tuning VLLMs to accomplish this task. However, fine-tuning VLLMs for specific tasks is often computationally expensive and risks catastrophic forgetting of general capabilities. To address this, we explore a generative approach that does not require additional fine-tuning, specifically leveraging a training-free paradigm using pre-trained VLLMs for table recognition. Recognizing the absence of dedicated benchmarks in this domain, we propose an evaluation benchmark based on a hierarchical design philosophy Sui et al. (2024); Cheng et al. (2025b); Liu et al. (2024b), comprising recognition tasks for table recognition. Through extensive evaluations, we identify a critical bottleneck: low-quality input images significantly hinder the table recognition capabilities of the evaluated VLLMs.
@@
To overcome this limitation, we propose the Neighbor-Guided Toolchain Reasoner (NGTR) framework for effective table recognition. One of the key features of the framework is its integration of lightweight models and the strategy of retrieval-augmented generation to improve image quality and guide structured data recognition. Specifically, we propose a preprocessing toolkit with various lightweight models to enhance input image quality. For each input instance, we retrieve a similar neighbor from the training data and use the experience gained from that neighbor to guide the generation of tool invocation plans. Furthermore, we incorporate a reflection-driven tool selection module at each step to iteratively refine the table recognition output. This enables VLLMs to produce more accurate structured data.
@@
To validate the effectiveness of the proposed NGTR framework, we conduct extensive experiments on multiple public table recognition datasets. The key observations are as follows: (1) Our NGTR framework significantly enhances the table recognition performance of naive VLLM-based approaches; (2) While VLLMs achieve competitive accuracy on specific datasets compared to traditional models, a noticeable performance gap remains in favor of traditional models. Nonetheless, we have preliminarily revealed the performance boundaries of VLLMs in several representative table recognition datasets. As is shown in Figure 1, the VLLM-based table recognition approach demonstrates the capability for universal modeling. This method facilitates a paradigm shift in design objectives from a model-centric to a data-centric focus, presenting significant potential for further exploration. We hope this work will inspire more research efforts in the future. In summary, the contributions of this paper are as follows:
  • We conduct a systematic investigation into VLLM-based table recognition by introducing a hierarchical benchmark for evaluating their recognition capabilities.
@@ 2 Related Work
Table Recognition. Earlier table recognition (TR) methods predominantly rely on heuristic rules Kieninger and Dengel (1999); Shigarov et al. (2016) or statistical learning techniques Ng et al. (1999), these approaches rely heavily on handcrafted features or implicit rules and show limited generalization ability. In the era of deep learning, numerous studies have made impressive progress in handling more intricate and heterogeneous table structures. Top-down methods Schreiber et al. (2017); Siddiqui et al. (2019); Ma et al. (2023); Qin et al. (2024) predict table borders to infer the structure information. Bottom-up methods Zheng et al. (2021); Qiao et al. (2021); Xing et al. (2023) first identify table cells with object detection models Ren et al. (2016); Carion et al. (2020), and then predict the cell relations to organize the row-column structures to form the overall tables. These methods follow an explicit two-stage learning paradigm with relatively strong transferability and explainability, yet the risks of ambiguous contents and boundless structures may lead to unstable and incorrect prediction results.
@@
Recently, sequence-based methods Zhong et al. (2020); Nassar et al. (2022); Huang et al. (2023) have been widely explored to directly generate markup sequence defining structures with specific decoders. Although these approaches require a massive of training data and computing resources, they have demonstrated substantial potential to unify visual-text parsing tasks Wan et al. (2024).
@@
Large Language Models. In recent years, LLMs have demonstrated exceptional performance in tasks such as multi-task learning Chen et al. (2024), zero-shot learning Kojima et al. (2022), and text generation Li et al. (2024). LLMs have not only broken through the limitations of traditional technologies in processing natural language text, but have also shown capabilities in reasoning Wei et al. (2022) and planning Guan et al. (2023); Gou et al. (2023). Meanwhile, VLLMs combine visual and language understanding capabilities, enabling LLMs to process visual information. For multimodal understanding scenarios (e.g., scene text recognition Wang et al. (2011), visual question answering Antol et al. (2015)), VLLMs have been widely validated as effective Guo et al. (2023); Ye et al. (2023); Liu et al. (2024a). With continuous progress in text-rich scenarios, some studies Zheng et al. (2024); Zhao et al. (2024); Chen et al. (2023) have also focused on enabling VLLMs to handle multimodal table understanding tasks.
@@
Despite their promising success in various domains, VLLMs applied to TR remain under-evaluated and under-explored. Our study presents a comprehensive benchmark for VLLM-based TR evaluation. Subsequently, we propose a novel framework to address the bottleneck of VLLMs, thereby enhancing their capabilities in TR.
3 Preliminary and Proposed Benchmark
3.1 Problem Definition
@@
3.2 Benchmark Evaluation Setup
@@
3.2.1 Recognition Task Design
Cell-level. We evaluate the cell-level recognition capability of VLLMs within three specific recognition tasks. Merged cell detection task aims to recognize cells that span multiple rows or columns in the table. Content/index-based cell recognition tasks evaluate the structural and content recognition capabilities of VLLMs, which are crucial for assessing whether VLLMs could perform well in fine-grained table recognition.
Row/Column-level. We evaluate the row/column-level recognition capability of VLLMs within two specific recognition tasks. Row/column index-based data recognition tasks are designed to assess whether VLLMs can accurately identify row/column elements.
Table-level. We evaluate the overall table-level recognition capability of VLLMs. The visual table size detection recognition task is designed to evaluate whether VLLMs can comprehend the global structural information and accurately determine the number of rows and columns in a table.
Table 1: Descriptions of the proposed hierarchical recognition tasks.
@@ 3.2.1 Recognition Task Design
3.2.2 Baselines
@@
In this study, we evaluate the performance of six VLLMs. For open-source VLLMs, we select Phi-3.5 (Phi) and Llama-3.2-90B (Llama) for evaluation. For closed-source VLLMs, we evaluat GPT-4o-mini (GPT-mini), Qwen-VL-Max (Qwen), GPT-4o (GPT) and Gemini-1.5-Pro (Gemini).
@@
Figure 2: Experimental results of VLLMs for the proposed hierarchical tasks. The tasks evaluated include the following: merged cell detection (MCD), content/index-based cell recognition (CCR, ICR), index-based row/column data recognition (IRDR, ICDR), and visual table size detection (VTSD).
3.3 Benchmark Evaluation Results
@@
Evaluation results of hierarchical recognition tasks are presented in Figure 2. Among all the VLLMs we selected, GPT and Gemini demonstrate the strongest performance, consistently outperforming the other VLLMs. Furthermore, the open-source Llama demonstrates a significant performance gap compared to the closed-source VLLMs. We give some highlights associated with the benchmark results as follows:
@@
Row-column Sensitivity Analysis. We found that all models show inconsistent performance between row/column-related tasks. To mitigate the influence of uneven distributions of rows and columns in the original data leading to varying difficulty levels, we further refine the experimental results by selecting samples where the difference between the number of rows and columns does not exceed three. The results are presented in Figure 3, showing that the accuracy for the column-related tasks is higher than for the row-related tasks, suggesting that VLLMs prefer to process column-structured data. This phenomenon likely occurs because columns usually exhibit high diversity, representing different attributes, whereas rows exhibit high similarity, typically corresponding to different entities. As a result, VLLMs are skilled in handling column-related tasks, where attribute information facilitates more accurate recognition.
Figure 3: Row-Column Sensitivity Analysis of VLLMs on Hierarchical Tasks with Gemini (Left) and GPT (Right).
@@
Image Quality Analysis. We conduct in-depth experiments and analysis on the TR task. Detailed experimental results are presented in Section 5. The results demonstrate that the VLLMs perform relatively well on the SciTSR dataset with higher image quality using simple prompts. However, the performance gap is considerably more significant when processing the PubTabNet dataset with lower image quality, especially compared to traditional models. This phenomenon indicates that the quality of the input image is a key bottleneck limiting the performance of VLLMs. Section 3.4 provides a more in-depth analysis of this bottleneck.
@@
Table 2: Specific description of the challenges and scenarios in low-quality image inputs.
@@
3.4 Bottleneck Analysis
We further investigate the performance of VLLMs under varying image quality conditions and assess their visual robustness to these conditions through empirical analysis.
3.4.1 Experimental Setup
To comprehensively evaluate the visual robustness of VLLM, we focus on three distinct visual challenges: the image quality challenge, the table border quality challenge, and the geometric deformation challenge. Details of these challenges are provided in Table 2.